# PBR-NeRF: Inverse Rendering with Physics-Based Neural Fields

### v1

@@ 4.2 Comparison with The State of The Art
Quantitative Results.  Table 1 shows that PBR-NeRF achieves state-of-the-art material estimation on the NeILF++ dataset, outperforming previous methods by 1–2 PSNR on average across albedo, metallicness, and roughness. We outperform all baselines in roughness and metallicness in most scenes across all illuminations. For albedo, we observe substantial gains over all competitors both on overall average and in Mix lighting scenarios, which are generally more complex due to their combined lighting setups. Despite performing slightly worse than NeILF [44] in novel view synthesis, we outperform the closely related NeILF++ [51] in novel view synthesis and all competitors in material estimation.
Among inverse rendering methods, PBR-NeRF achieves state-of-the-art results in novel view synthesis on DTU, as shown in Table 2. We report the average results over three training runs for both our approach and NeILF++ [51], mitigating the impact of training variability. Our method achieves an improvement of 0.37 in mean PSNR over NeILF++ and substantially outperforms all other methods. Additionally, PBR-NeRF ranks first in 10 scenes, with PSNR gains ranging from 0.12 to 2.51 over NeILF++, demonstrating consistent improvements over our baseline.
DTU Scan 37 DTU Scan 40 DTU Scan 63
RGB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Roughness Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Metallicness Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeILF++ [51] Ours NeILF++ [51] Ours NeILF++ [51] Ours
@@
Qualitative Results.  We further validate our material estimation performance qualitatively against NeILF++ on the NeILF++ dataset in Fig. 4 (see Appendix Sec. B for additional qualitiative results). Over all three challenging Mix lighting conditions, our physics-based losses improve both lighting and material estimation. Our estimated lighting exhibits lower entropy with more concentrated sources and fewer artifacts, especially in the upper halves of environment maps. We also eliminate fringes and patches generated on the sphere and cube by NeILF++ due to “baked-in” specular highlights and achieve effective separation of diffuse and specular BRDF components via our NDF-weighted Specular Loss. Most importantly, we consistently outperform NeILF++ on albedo estimation, which is a critical challenge in inverse rendering. Note that although NeILF++ surpasses our metallicness estimation on Studio Mix, it fails significantly to recover the correct albedo. In general, our predicted albedo is—correctly—less bright than for NeILF++, with the floor, cube, and sphere all free from incorrect specular highlights and artifacts.
We also compare our material estimation on DTU in Fig. 5. On scan 37, we improve metallicness and roughness, accurately predicting the scissor and clamp as metal, unlike NeILF++, which creates artifacts near the scissor handle. NeILF++ also predicts the scissor and part of the clamp as rough material instead of metal. This issue creates patchy artifacts both in the metallicness and roughness. On scan 40 with the rough bricks, our roughness estimate on shadows is better but we also exhibit albedo artifacts. On scan 63, we successfully recover fine details such as the green apple stem, which NeILF++ misses in its material predictions.
@@ 4.3 Ablation Study
We perform an ablation study of the proposed physically-based losses in Tab. 3 to individually evaluate each loss’s contribution. A qualitative analysis is provided in Appendix Sec. A. The baseline (ID 1) is NeILF++ [51] without the Lambertian loss. Adding the Conservation of Energy Loss (ID 2) improves material estimation, with albedo PSNR increasing by 0.84 and metallicness PSNR by 0.8, respectively. In our final method (ID 3), adding the specular loss $\mathcal{L}_{\text{spec}}$ further improves the albedo and metallicness. Specifically, the albedo PSNR increases by 2.02 PSNR while the albedo SSIM improves by 6.99%. Although the metallicness PSNR drops by 0.10 PSNR from ID 2, the metallicness SSIM increases by 9.45%. Overall, our complete method (ID 3) with both physics-based losses improves metallicness by 0.70 PSNR and albedo by 2.86 PSNR over the baseline (ID 1), while maintaining comparable performance in RGB and roughness.
Table 3: Ablation of our physics-based losses. Performance of ablated versions of PBR-NeRF and of its full version is compared on the City illumination using PSNR and SSIM for Disney BRDF parameters and novel views.
@@ 5 Conclusion
We present PBR-NeRF, a novel inverse rendering method leveraging neural fields optimized with PBR-inspired losses. These losses provide physically valid inductive biases, enabling the neural fields to better disambiguate materials from lighting—addressing a major challenge in inverse rendering. Thorough experiments across various benchmarks demonstrate the effectiveness of PBR-NeRF in not only material estimation and inverse rendering, but also in novel view synthesis. In particular, we qualitatively show a compelling disentanglement of the diffuse and specular components in our estimated BRDFs. We believe that our physics-based approach will inspire further work on PBR-driven, decomposed neural fields.
@@ Appendix A Ablation Study Qualitative Results
We present qualitative results in Fig. 6 for the ablation study of the proposed physics-based losses on the NeILF++ dataset City scene in Tab. 3 of Sec. 4.3.
To ensure a fair comparison, all ablated versions share the same hyperparameters, varying only the inclusion of the proposed loss terms. For consistency with NeILF++ [51], we use a lower number of sampled incident light directions $|S_{L}|=128$ during training instead of our final method’s $|S_{L}|=256$ to allow direct comparison with our reproduced NeILF++ [51] results on the City scene in Tab. 1.
@@
Adding the Conservation of Energy Loss (ID 2) mitigates these artifacts by enforcing energy conservation for the diffuse and specular lobes. As a result, the lighting estimation is also improved since the estimated light does not need to compensate for the previously non-energy-conserving (overly reflective) predicted materials. The predicted lighting is more consistent and improves the originally overly dim patches in the bottom left corner of the baseline (corresponding to the cube). Reducing specular artifacts in the lighting also improves the material estimation, as shown by the increased albedo and metallicness PSNR in Tab. 3. Qualitatively, we also see that the albedo of the helmet and cube are effectively recovered in all regions except the shadowed area. Despite these improvements, there are still fringing artifacts on the sphere as well as shadowing effects on the ground and cube.
Incorporating the specular loss $\mathcal{L}_{\text{spec}}$ in our final method (ID 3) improves the albedo and metallicness even further. We once again highlight that using both our novel physics-based losses improves metallicness by 0.7 PSNR and albedo by 2.86 PSNR over the baseline (ID 1). Qualitatively, the albedo estimation, which is most important, is significantly improved in our final method (ID 3), especially with the green albedo of the helmet and the gray albedo of the cube. However, there are still limitations with our proposed method, which highlights directions for future improvement. In particular, we fail to fully reproduce the specular effects in the helmet’s reflective visor and metallic pieces in the predicted RGB. This limitation is likely due to our overestimation of roughness and the missing fine details in our metallicness prediction. Furthermore, we still exhibit limited specular effects in our predicted albedo with fringing on the sphere and shadowy patches on the cube and floor.
@@
RGB Refer to caption Refer to caption Refer to caption Refer to caption
Roughness Refer to caption Refer to caption Refer to caption Refer to caption
Metallicness Refer to caption Refer to caption Refer to caption Refer to caption
Albedo Refer to caption Refer to caption Refer to caption Refer to caption
Baseline (ID 1) $\mathcal{L}_{\text{cons}}$ (ID 2) $\mathcal{L}_{\text{cons}}$, $\mathcal{L}_{\text{spec}}$ (ID 3) Ground Truth
@@ Appendix B Additional Qualitative Results
Appendix C Additional Background Details
For completeness, we provide further implementation details on the NeILF++ implicit differential renderer (Sec. C.1), joint material-illumination-geometry optimization (Sec. C.2), and hyperparameter sweeps (Sec. C.3).
C.1 Implicit Differential Renderer (IDR)
@@ C.2 Joint Optimization
C.3 Hyperparameter Sweeps
We now specify how we picked various hyperparameters introduced with our novel Conservation of Energy Loss (7) and NDF-weighted Specular Loss (8).

### v2

@@ 4.2 Comparison with The State of The Art
Quantitative Results.  Table 1 shows that PBR-NeRF achieves state-of-the-art material estimation on the NeILF++ dataset, outperforming previous methods by 1–3 PSNR on average across albedo, metallicness, and roughness. We outperform all baselines in roughness and metallicness PSNR in most scenes across all illuminations. For albedo, we observe substantial gains in PSNR and SSIM over all competitors both on overall average and in Mix lighting scenarios, which are generally more complex due to their combined lighting setups. Our closest competitor for albedo estimation is SG-Env, which combines the PhySG [53] Spherical Gaussian (SG) lighting representation with a spatially-varying microfacet BRDF. SG-Env estimates albedo better on the Env scenes since its explicit SG lighting models environment lighting better than PBR-NeRF’s under-constrained implicit lighting MLP. However, PBR-NeRF performs over 3 albedo PSNR better than SG-Env on the more challenging Mix scenes, as SGs cannot model spatially varying illumination such as mixed light sources, indirect lighting, and occlusions. While SG-Env struggles with mixed, SV lighting, PBR-NeRF estimates albedo equally well across environment and mixed lighting. Despite performing slightly worse than NeILF [44] and NeILF++ [51] in metallicness SSIM, we outperform all competitors in all other material estimation metrics.
Among inverse rendering methods, PBR-NeRF achieves state-of-the-art results in novel view synthesis and geometry estimation on DTU, as shown in Table 2. We report the average results and standard deviation over three training runs both for our approach and NeILF++ [51], mitigating the impact of training variability. Our method achieves an improvement of 0.78 in mean PSNR and 0.339 mm in Chamfer Distance over NeILF++ with lower variance, showing that our proposed losses positively impact geometry estimation too. Additionally, PBR-NeRF ranks first in 14 scenes for RGB PNSR, with PSNR gains ranging from 0.12 to 2.56 over NeILF++, demonstrating consistent improvements over our closest baseline, and it substantially outperforms all other methods.
DTU Scan 37 DTU Scan 63 DTU Scan 105
RGB
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Normals
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Roughness
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Metallicness
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference NeILF++ [51] Ours Reference NeILF++ [51] Ours Reference NeILF++ [51] Ours
@@
Qualitative Results.  We further validate our material estimation performance qualitatively against NeILF++ on the NeILF++ dataset in Fig. 4 (see supplement for additional qualitative results). Over all three challenging Mix lighting conditions, our physics-based losses improve both lighting and material estimation. Our estimated lighting exhibits lower entropy with more concentrated sources and fewer artifacts, especially in the upper halves of environment maps. We also eliminate fringes and patches generated on the sphere and cube by NeILF++ due to “baked-in” specular highlights and achieve effective separation of diffuse and specular BRDF components via our NDF-weighted Specular Loss. Most importantly, we consistently outperform NeILF++ on albedo estimation, which is a critical challenge in inverse rendering. In particular, our albedo estimation in Studio Mix properly recovers the helmet and visor whereas NeILF++ fails to recover any helmet details and incorrectly predicts a black albedo. In general, our predicted albedo is—correctly—less bright than for NeILF++, with the floor, cube, and sphere all free from incorrect specular highlights and artifacts.
We compare our DTU material and geometry estimation in Fig. 5. On scan 37, we improve metallicness and roughness, accurately predicting the scissor and clamp as metal, unlike NeILF++, which creates artifacts near the scissor handle in its predicted normals, metallicness, and roughness. On scan 63, we successfully recover fine details like the green apple stem, which NeILF++ misses in its predicted normals and materials. On scan 105, we correctly reproduce the brick’s top surface, whereas NeILF++ has RGB, normals, and albedo artifacts. Our roughness and metallicness on the orange toy fur are also more consistent than NeILF++ for the homogeneous furry material.
@@ 4.3 Ablation Study
We perform an ablation study of the proposed physically-based losses in Tab. 3 to individually evaluate each loss’s contribution. A qualitative analysis is provided in the supplement. The baseline (ID 1) is NeILF++ [51] without the Lambertian loss. Adding only the Conservation of Energy Loss $\mathcal{L}_{\text{cons}}$ (ID 2) improves material estimation, with albedo PSNR increasing by 2.34 and roughness PSNR by 0.28, respectively. Adding only the specular loss $\mathcal{L}_{\text{spec}}$ (ID 3) also improves material estimation over ID 1, with albedo PSNR increasing by 2.35, roughness PSNR by 0.25, and metallicness PSNR by 0.26, respectively. In our final method (ID 4), adding both $\mathcal{L}_{\text{cons}}$ and $\mathcal{L}_{\text{spec}}$ achieves the highest albedo, roughness, and metallicness PSNR. Specifically, albedo PSNR and SSIM increase by an additional 0.93 PSNR and 1.84% over ID 3, respectively. Although metallicness SSIM is lower for ID 2–4 compared to ID 1, metallicness PSNR is improved. Overall, our complete method (ID 4) with both physics-based losses improves roughness by 0.35 PSNR, metallicness by 0.27 PSNR, and albedo by 3.28 PSNR over the baseline (ID 1), while maintaining on-par performance in RGB.
Table 3: Ablation of our physics-based losses. Performance of ablated versions of PBR-NeRF and of its full version is compared on the mean PSNR and SSIM of all 6 scenes of the NeILF++ dataset  [51] for Disney BRDF parameters and novel views.
@@ 5 Conclusion
We present PBR-NeRF, a novel inverse rendering method leveraging neural fields optimized with PBR-inspired losses. These losses provide physically valid inductive biases, enabling our neural fields to better disambiguate materials from lighting—addressing a major challenge in inverse rendering. Thorough experiments across various benchmarks demonstrate the effectiveness of PBR-NeRF not only in material estimation and inverse rendering, but also in novel view synthesis. In particular, we qualitatively show a compelling disentanglement of the diffuse and specular components in our estimated BRDFs. We believe that our physics-based approach will inspire further work on PBR-driven, decomposed neural fields.
@@ Appendix A Ablation Study Qualitative Results
We present qualitative results in Fig. 6 for the ablation study of the proposed physics-based losses on the NeILF++ dataset City scene in Tab. 3 of Sec. 4.3. To ensure a fair comparison, all ablated versions share the same hyperparameters as our proposed method described in Sec. 4.1, varying only the inclusion of the proposed loss terms.
@@
Adding the Conservation of Energy loss (ID 2) mitigates these artifacts by enforcing energy conservation for the diffuse and specular lobes. As a result, the lighting estimation is also improved since the estimated light does not need to compensate for the previously non-energy-conserving (overly reflective) predicted materials. The predicted lighting is more consistent and improves the originally overly dim patches in the bottom left corner of the baseline (corresponding to the cube). Reducing specular artifacts in the lighting also improves the material estimation, as shown by the increased albedo, roughness, and metallicness PSNR in Tab. 3. Qualitatively, we also see that the albedo of the helmet and cube are effectively recovered in all regions except the shadowed area. Despite these improvements, there are still fringing artifacts on the sphere as well as shadowing effects on the ground and cube.
Adding the specular loss instead of the Conservation of Energy loss (ID 3), shows similar improvements in albedo, roughness, and metallicness over the baseline (ID 1).
Finally, incorporating both the Conservation of Energy and specular losses in our final method (ID 4) improves the albedo, roughness, and metallicness even further. We once again highlight that using both our novel physics-based losses improves roughness by 0.35 PSNR, metallicness by 0.27 PSNR, and albedo by 3.28 PSNR over the baseline (ID 1). Qualitatively, the albedo estimation, which is most important, is significantly improved in our final method (ID 4), especially with the green albedo of the helmet and the gray albedo of the cube. The roughness and metallicness of the shadowed area of the cube are also recovered with fewer artifacts. However, there are still limitations with our proposed method, which highlights directions for future improvement. In particular, we fail to fully reproduce the specular effects in the helmet’s reflective visor and metallic pieces in the predicted RGB. This limitation is likely due to our overestimation of roughness and the missing fine details in our metallicness prediction. Furthermore, we still exhibit limited “baked-in” specular effects in our predicted albedo with fringing on the sphere and incorrectly predict shadowy patches on the cube and floor.
@@
RGB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Roughness Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Metallicness Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Baseline (ID 1) $\mathcal{L}_{\text{cons}}$ (ID 2) $\mathcal{L}_{\text{spec}}$ (ID 3) $\mathcal{L}_{\text{spec}}$ (ID 4) Ground Truth
@@ Appendix B Additional Qualitative Results
Appendix C Non-Energy Conserving BRDFs
We extend the discussion in Section 3.2 and provide additional intuition for when the Conservation of Energy property for BRDFs fails to hold. The Disney BRDF creates excess energy for unnormalized terms or when the Schlick Fresnel approximation is invalid, causing overly bright specular highlights [10]. Further, energy can be lost as the underlying microfacet BRDF models rough materials with single reflection instead of multiple scattering on the microsurface.
Appendix D Additional Background Details
For completeness, we provide further implementation details on the NeILF++ implicit differential renderer (Sec. D.1), joint material-illumination-geometry optimization (Sec. D.2), and hyperparameter sweeps (Sec. D.3).
D.1 Implicit Differential Renderer (IDR)
@@ D.2 Joint Optimization
Note that the NeRF SDF is used twice per training sample during the Joint phase: (1) estimating $L_{\text{o,NeRF}}$ for $\mathcal{L}_{\text{geo}}$ and $\mathcal{L}_{\text{ref}}$; (2) sphere tracing to compute $\mathbf{x}$ and $\mathbf{n}$ for $L_{\text{o,PBR}}$.
D.3 Hyperparameter Sweeps
We now specify how we picked various hyperparameters introduced with our novel Conservation of Energy Loss (7) and NDF-weighted Specular Loss (3.2).